I need help with a solve by substitution method equation< the answer should be an ordered pair.
x= 24 - 4y
4x + 3y = -8
Any help would be greatly appreciated!!!
Blaine

Answer by ewatrrr(24785)


Hi
Solving by substitution method
x = 24 - 4y
4x + 3y = -8
4(24-4y) + 3y = -8
96 - 16y + 3y = -8
104 = 13y
8 = y & x = 24-4y = -8 Ordered pair (-8,8)is the solution
CHECKING our Answer***
4x + 3y = -8
-32 + 24 = -8